How many moles of oxygen gas (O2) are required to react with 0.67 moles of iron (Fe)


1
Expert's answer
2021-02-03T07:24:05-0500

3Fe + 2O₂ = Fe₃O₄

"\nu" (O2)="\nu" (Fe)*2/3= 0.67*2/3=0.45 mol